OpenStudy (anonymous):

What are the solutions of 6x2 − x − 2 = 0?

OpenStudy (mathstudent55):

You can try factoring, or you can use the quadratic formula.

OpenStudy (anonymous):

6x^2 - x - 2 = 0 Factor the equation. (3x - 2)(2x + 1) = 0 Set each binomial equal to zero. 3x - 2 = 0 and 2x + 1 = 0 Solve both equations. x = 2/3 or x = - 1/2 Check answers(optional). 6(- 1/2)^2 - (- 1/2) - 2 = 0 6(1/4) + 1/2 - 2 = 0 6/4 + 1/2 - 2 = 0 6/4 + 2/4 - 8/4 = 0 0 = 0 6(2/3)^2 - (2/3) - 2 = 0 6(4/9) - 2/3 - 2 = 0 24/9 - 2/3 - 2 = 0 24/9 - 6/9 - 18/9 = 0 0 = 0

OpenStudy (anonymous):

well thank you that. looked like it took 4ever